OpenStudy (anonymous):

The cost, c in dollars of car rental is 8 + .25m, where m is the number of miles driven. Use the equation to estimate the cost of car rental if the number of miles driven is 33. A. about 35 dollars B. about 12 dollars C. about 21 dollars D. about 16 dollars

OpenStudy (king):

about 16 dollars c=8+0.25m c=8+0.25*33 c=8+8.25 c=16.25

OpenStudy (anonymous):

thank you. I dont know how you know all this. Its confusing as hell

OpenStudy (anonymous):

just read the question well twice then analyse wat is given then u wud figure out that m=33 given so just plug it in take time to write down the given data first

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

vince what level are you at academically...

OpenStudy (anonymous):

well this is the last class i need for my BA in sports and health sciences